OpenStudy (anonymous):

What is square root of 108 in simplified radical form?

OpenStudy (anonymous):

square root of 108 meaning \[\sqrt{108}\]

OpenStudy (igreen):

Break it into factors.

OpenStudy (igreen):

108 Find the factors..find one that is a perfect square.

OpenStudy (anonymous):

A. \[\sqrt[27]{4}\] B. \[\sqrt[4]{27}\] C. \[\sqrt[6]{3}\] D. \[\sqrt[3]{6}\]

OpenStudy (igreen):

3 and 36.

OpenStudy (anonymous):

I think C

OpenStudy (anonymous):

am i correct?

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

nooo

OpenStudy (anonymous):

and here u goo correct

OpenStudy (anonymous):

Lol thanks ^_^